# WIMP/FIMP dark matter and primordial black holes with memory burden effect

### v1

@@ 4 Summary
The lifetime of PBHs has recently been discovered to be extended by the memory burden effect. Hence, light PBHs may currently survive and become candidates for DM. We assume that DM is made of thermally produced WIMPs (or FIMPs), WIMPs (or FIMPs) produced via Hawking radiation of PBHs, and survived PBHs via the memory burden effect. If the effect of thermal production of WIMPs (or FIMPs) is much smaller than that of particle production of PBHs via Hawking radiation, we can omit the contribution of thermally produced WIMPs (or FIMPs), obtaining the case already studied in [23]. Thus, we focus on the case that the effect of thermal production of WIMPs (or FIMPs) is much larger than that of particle production of PBHs via Hawking radiation.
Results from numerical calculations show that the relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect, $q\lesssim\mathcal{O}(10^{-3})$, is only allowed for $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$ if DM is made of thermal WIMPs, WIMPs from PBHs, and survived PBHs. A similar result is obtained if we consider FIMPs instead of WIMPs as the DM components.
Although the results of this study can be inferred through qualitative analyses, we believe that performing a quantitative analysis to confirm the qualitative prediction is important for increasing the accuracy and significance of the findings.

### v3

@@ 5 Summary
The lifetime of PBHs is extended by the memory burden effect. Hence, light PBHs may survive and become candidates for DM. We assume that DM is made of thermally produced WIMPs/FIMPs, WIMPs/FIMPs produced via the Hawking radiation of PBHs, and PBHs that survived via the memory burden effect. We can omit the contribution of thermally produced WIMPs/FIMPs to DM relic abundance, obtaining the case studied by [25], if the effect of the thermal production of WIMPs/FIMPs is much smaller than that of PBH particle production via Hawking radiation. Thus, we focus on a case where the effect of the thermal production of WIMPs/FIMPs substantially exceeds that of PBH particle production via Hawking radiation.
The relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect, $q\lesssim\mathcal{O}(10^{-3})$, is only allowed at $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$ for DM consisting of thermal WIMPs, WIMPs from PBHs, and survived PBHs. A similar result is obtained if we consider FIMPs instead of WIMPs as the DM components.
We also find that the WDM constraint–the lower limit of the DM mass, which is related to structure formation in the universe–decreases with the memory burden effect of PBHs. The WDM constraint is more effective for FIMPs than for WIMPs.